# The Measurement Imbalance in Agentic AI Evaluation Undermines Industry Productivity Claims

### v1

@@ 1 Introduction
By grounding what we count as evidence, this paper aims to reset the conversation on agentic AI efficiency, not to constrain innovation, but to ensure that the claims we make reflect the systems we actually build (see Section A.4 for formal definitions).
@@ 2 Why Evaluation Matters
MLAgentBench huang_mlagentbench_2024 is an example of today’s technical focus: agents are scored on Pass@k success, token efficiency, and wall-clock time while running scripted ML experiments. The benchmark yields crisp, automatable numbers but tells us nothing about how well an agent fits a collaborative workflow or whether its behavior degrades after extended autonomous operation.
By contrast, TrAAIT, a clinician-trust instrument for AI decision support, captures perceived usefulness, ease of integration, and social influence through validated survey scales stevens_theory_2023 . TrAAIT’s scores directly predict real-world adoption but is rarely included in research papers. The field celebrates technical benchmark leaderboards while few report trust or usability scores, reflecting a marked misalignment between what we measure and what determines deployment success.
Systems that ace technical benchmarks can still be rejected by end-users if trust, workflow compatibility, or explanation quality are weak (adoption risk). Point-in-time accuracy masks drift, emergent behavior, and edge-case failure modes that surface only after deployment, creating potential for harm that remains invisible to current evaluation frameworks (safety blind spots). ROI projections based solely on speed or accuracy overlook human oversight costs and adaptation lags, leading to costly rollout reversals and failed investments (misjudged productivity).
@@ 3.1 Methodology
We conducted systematic searches across arXiv and Google Scholar. Search terms included combinations of “agentic AI”, “AI agents”, “LLM agents”, “evaluation”, “benchmark”, “assessment”, and “metric”. Papers were included if (1) published between January 2023 and April 2025, (2) focused on agentic or LLM-based systems as defined in Section 1, (3) available in English, (4) proposed or applied evaluation methodology and metrics.
Two reviewers independently screened 138 papers (with 85% agreement), with disagreements resolved by a third reviewer. Out of 138 papers, 84 of them met our inclusion criteria. We adopted a four-category scheme that was developed through an iterative process based on established AI evaluation metrics xia2024ai , pilot coding of 25 papers, and inter-reliability testing. These four metric categories are: Technical such as task success rate, accuracy, latency, throughput, human-centered such as trust, usability, collaboration, workflow integration, safety and governance such as robustness, compliance, explainability, alignment, and economic impact such as ROI, cost savings, value realization. We also looked at evaluation quality to make sure the methodology details are described in the paper. Coding was binary per category. The full codebook is available in Section A.1, and coded records are available in the accompanying PDF file.
@@ 3.2 Data Analysis
In our analysis of qualified papers technical metrics dominated (83%), while human-centered and economic impact metrics were less common (both 30%). Notably, only 15% of papers included both technical and human-centered metrics, and a mere 5% incorporated any longitudinal dimension (detailed distributions in Section A.3).
@@ 5 A Balanced Framework for Evaluating Agentic AI
The failures highlighted in Section 4 are not isolated accidents. They reflect a systemic overemphasis on technical correctness at the expense of human, temporal, and contextual factors. To realign evaluation with real-world success, we propose a framework that balances these dimensions, supports domain adaptation, and maintains practical feasibility.
@@ 5.2 Dimensional Interdependence
A key insight from our case studies is that dimensions don’t operate in isolation. Instead, they form a complex, interdependent system where changes in one dimension inevitably affect others (Fig. 1).
@@ 5.4 Metric Interaction Formalism
In current practice, $w_{T}\approx 1$ and all others are near zero, implicitly treating technical success as a proxy for overall value. But evidence from Section 3 and Section 4 shows this fails to predict actual outcomes.
@@
Organizations can implement this framework through a phased approach that balances thoroughness with practical constraints.Section A.2 provides a practical implementation example.
@@ A.1 Coding Schema for Meta-Analysis
Each paper was coded using a binary vector [T, H, S, E], corresponding to Table 1.
@@ A.2 Practical Implementation Process
This phased approach balances comprehensiveness with real-world constraints, allowing organizations to begin shifting toward balanced evaluation without disrupting innovation cycles (Fig. 2).
@@ A.3 Metric Type Distribution
Fig. 3 displays three measures for each metric type: absolute number of papers, percentage of all examined papers, and percentage of qualified papers that met quality thresholds. Note that individual papers often employed multiple evaluation approaches.

### v3

@@ 1 Introduction
By grounding what we count as evidence, this paper aims to reset the conversation on agentic AI efficiency, not to constrain innovation, but to ensure that the claims we make reflect the systems we actually build (see Section˜A.4 for formal definitions).
@@ 2 Why Evaluation Matters
MLAgentBench huang_mlagentbench_2024 is an example of today’s technical focus: agents are scored on Pass@k success, token efficiency, and wall‑clock time while running scripted ML experiments. The benchmark yields crisp, automatable numbers but tells us nothing about how well an agent fits a collaborative workflow or whether its behavior degrades after extended autonomous operation.
By contrast, TrAAIT, a clinician‑trust instrument for AI decision support, captures perceived usefulness, ease of integration, and social influence through validated survey scales stevens_theory_2023 . TrAAIT’s scores directly predict real-world adoption but is rarely included in research papers. The field celebrates technical benchmark leaderboards while few report trust or usability scores, reflecting a marked misalignment between what we measure and what determines deployment success.
Systems that ace technical benchmarks can still be rejected by end‑users if trust, workflow compatibility, or explanation quality are weak (adoption risk). Point‑in‑time accuracy masks drift, emergent behavior, and edge‑case failure modes that surface only after deployment, creating potential for harm that remains invisible to current evaluation frameworks (safety blind spots). ROI projections based solely on speed or accuracy overlook human oversight costs and adaptation lags, leading to costly rollout reversals and failed investments (misjudged productivity).
@@ 3.1 Methodology
We conducted systematic searches across arXiv and Google Scholar. Search terms included combinations of “agentic AI”, “AI agents”, “LLM agents”, “evaluation”, “benchmark”, “assessment”, and “metric”. Papers were included if (1) published between January 2023 and April 2025, (2) focused on agentic or LLM-based systems as defined in Section˜1, (3) available in English, (4) proposed or applied evaluation methodology and metrics.
Two reviewers independently screened 138 papers (with 85% agreement), with disagreements resolved by a third reviewer. Out of 138 papers, 84 of them met our inclusion criteria. We adopted a four-category scheme that was developed through an iterative process based on established AI evaluation metrics xia2024ai , pilot coding of 25 papers, and inter-reliability testing. These four metric categories are: Technical such as task success rate, accuracy, latency, throughput, human-centered such as trust, usability, collaboration, workflow integration, safety and governance such as robustness, compliance, explainability, alignment, and economic impact such as ROI, cost savings, value realization. We also looked at evaluation quality to make sure the methodology details are described in the paper. Coding was binary per category. The full codebook is available in Section˜A.1, and coded records are available in the accompanying PDF file.
@@ 3.2 Data Analysis
In our analysis of qualified papers technical metrics dominated (83%), while human-centered and economic impact metrics were less common (both 30%). Notably, only 15% of papers included both technical and human-centered metrics, and a mere 5% incorporated any longitudinal dimension (detailed distributions in Section˜A.3).
@@ 5 A Balanced Framework for Evaluating Agentic AI
The failures highlighted in Section˜4 are not isolated accidents. They reflect a systemic overemphasis on technical correctness at the expense of human, temporal, and contextual factors. To realign evaluation with real-world success, we propose a framework that balances these dimensions, supports domain adaptation, and maintains practical feasibility.
@@ 5.2 Dimensional Interdependence
A key insight from our case studies is that dimensions don’t operate in isolation. Instead, they form a complex, interdependent system where changes in one dimension inevitably affect others (Fig.˜1).
@@ 5.4 Metric Interaction Formalism
In current practice, $w_{T}\approx 1$ and all others are near zero, implicitly treating technical success as a proxy for overall value. But evidence from Section˜3 and Section˜4 shows this fails to predict actual outcomes.
@@
Organizations can implement this framework through a phased approach that balances thoroughness with practical constraints.Section˜A.2 provides a practical implementation example.
@@ A.1 Coding Schema for Meta-Analysis
Each paper was coded using a binary vector [T, H, S, E], corresponding to Table˜1.
@@ A.2 Practical Implementation Process
This phased approach balances comprehensiveness with real-world constraints, allowing organizations to begin shifting toward balanced evaluation without disrupting innovation cycles (Fig.˜2).
@@ A.3 Metric Type Distribution
Fig.˜3 displays three measures for each metric type: absolute number of papers, percentage of all examined papers, and percentage of qualified papers that met quality thresholds. Note that individual papers often employed multiple evaluation approaches.